Thursday, November 27, 2014

Marantz 2325 Receiver

Marantz 2325 Receiver
Got this unit in on a commission, the owner had replaced the large lyrics prior to coming onto my bench
So far this has got to be right up there with the 2230 as a favorite unit, clean and easy to work on. Very well built and sounds just amazing, with a modern clarity to it.

P800 Power Supply/Relay
The first thing I noted the lytics had signs of obvious heat exposure present with the shrinking and the large capacitor at C803 had leaked lytic material all over the near by resistors including R816 in the control network. At some point someone had replaced the Omron relay with a MY-2 type, though an LY is standard for the 2325.

All the electrolytic were replaced with high temp Panasonic FC's which seem to do quite well in the Psupply. Relay protection diodes H808/H809 131B  were replaced with like to-220 MUR162 and H810 WZ-140 diode with a 1N5422. The coated resistors were replaced with 5% carbon films at rated value. The relay was replaced with the specified Omron LY-2 24VDC type. All the Glue was soaked with acetone and gently removed. The regulated transistors were regressed. Voltage adjustments at R809 were made to 35.04VDC.
Before Re-work


The amplifier board was all stock and untouched, a lot of debate went on about films, the greens and browns I like to keep, but the coupling capacitor C701 had several options, so far its still original, but stack, or a PP would be excellent as were suggested to me. All the electrolytic were replaced with Nichicon FG's and Silmic II's at C707/706/C703. C710/C711 ceramics in the feedback stage and C713 were replaced with C0G/NPO's type. A service bulletin was performed replacing H718/719 VD1122 diodes with single 1N4148's. All the TO-3 mica insulators were replaced and regreased,a long with the top mounted diode. Offset was monitored and set at 0.001-VDC left and right and idle current adjusted to 20mv across. Initially the bias pots R740 were damn near stuck in place but with some faithful D5 and some action I got it too adjust properly. I want to thank Ben here for the simple suggestion going CCW and bring it up DBT at half the VAC, good trick.

As with most of the other boards it was stock as well. We have several diodes to update here also. The axial was replaced with a 220uf Nichicon axial, with Silmic IIs at C403/404. Tantalums C401/402 were replaced with  ECQ stacked film.
Varistor H409 a VD1212 was replaced with two 1N4148 diodes in series and heat-shrinked and insulating the legs.  H408/407  2473's with a single 1N4148. The prone to noise H405/406 2SC458's were replaced with the common KSC1845 transistor.

All the electrolytic were replaced with Silmic II's and Nichicon FG's. CE37/38 were replaced with bipolar MUSE ES capacitors. The tantalums were replaced with polar FG's. Signal ceramics were replaced with C0G/NPO at CE07/08. CE05/06 where left at this time, though some PP could be used here I think with nice results. All the Pots were cleaned and treated at this time.

Where signals are converted and passed to the dolby. All the electrolytic were replaced with Silmic II's and a Nichicon FG at CH13. The varistor VD1222 at HH11 was replaced with two  1N4148 diodes in series and heat-shrinked, insulating the legs. PL01 the Dolby, tone and meter board, the electrolytic were replaced with Silmic II's.  The meter capacitors were also replaced with same value Silmics.

No comments:

Post a Comment